Compute the charge of the following ions. Show your solution.


1. Anion with a charge of -3



2. Cation with a charge of +2



3. Anion with a charge of +5



1
Expert's answer
2022-02-11T08:47:21-0500


The net charge of the following ions:

1. Anion with a charge of -3


"Q=-3e=-3*1.6*10^{-19}\: \rm C=-4.8*10^{-19}\: \rm C"

2. Cation with a charge of +2


"Q=2e=2*1.6*10^{-19}\: \rm C=3.2*10^{-19}\: \rm C"


3. Anion with a charge of +5

"Q=5e=5*1.6*10^{-19}\: \rm C=8.0*10^{-19}\: \rm C"